If the length L of a rectangle is 3 meters more than twice its width and its perimeter is 300 meters. How do you solve the equation?
Answer by stanbon(75887)

If the length L of a rectangle is 3 meters more than twice its width and its perimeter is 300 meters. How do you solve the equation?
-----------
L = 2W+3
W is the width
---------------------
Perimeter = 2(L+W)
300 = 2(2W+3 + W)
150 = 3W+3
3W = 147
W = 49 meters (width)
L = 2W+3 = 98+3 = 102 meters (length)
=============
Cheers,
Stan H.
============